LBTO Technical/Science Night

OSA: Steve Allanson

ISp: Jenny Power

ISA: B. Rothberg (Tucson)

Manager: John Little

=========================================================================================

UT 12:31 - LBC up and running v2.1

UT 01:07 - Opening Enclosure UT 01:30 - Sending Skyflat U and TiO test OB

UT 01:38 - Taking U-Bessell and TiO Skyflats -scale 2x (PA=0)

UT 01:42 - U Bessell range from 11k-8k, TiO go from 10k - 3k - running 1 more time

UT 01:43 - scaling by 7x U and TiO (PA=0)

UT 01:46 - TiO ~5k-4k (10 exposures in these counts should be sufficient S/N) - U-Bessell looks good - use these for LBTO flats

UT 01:49 - Taking B-Bessell and CN Skyflats PA=0 - scaling by 20x

UT 01:50 - CN ~4-5k first image, 2k last image

UT 01:53 - B-Bessell 30-40k on first image (too high)

UT 01:57 - Rerunning B-CN flats PA=0 scale 45x

UT 02:00 - CN ~2k in first image - stopping after 2 exposures in CN and 1 in B-Bessell

UT 02:03 - Attempting V and I scale by 30x - stopping after two exposures - too dark and too many stars visible

==========================================================================================

UT 02:10 - sending copointing OB for DOFPIA

UT 02:23 - DOFPIA 1.1" Red 1.8" blue, but we are at airmass 1.8 - will copoint now and then wait a bit for lower airmass

UT 02:31 - Rerunning DOFPIA - airmass 1.74

==========================================================================================

UT 02:34 - M31substr6.OB - airmass 1.72, seeing ~1.1" Blue; 0.8" Red

UT 02:38 - DIMM reports 0.59"

UT 02:45 - V-Bessell shows a lot of saturated stars in long exposure

UT 02:57 - Zero counts on the red tec chips ~20min - IT #3759

UT 02:59 - Stopping M31substr6.OB - red side not guiding- turning LBC ON/OFF again - completed 8/18 exposures in OB

UT 03:00 - first zero 02:41 UT - three LBC-R images look OK, but call was to stop now, instead of continuing without focus and wavefront sensor correction

UT 03:06 - LBC back on, Running DOFPIA since its been 40min since last one and now at airmass 1.50

UT 03:09 - DOFPIA converged, restarting M31substr6.OB (only 1 dther position completed in original OB, and part of 2nd done, instead of hacking together and wasting time creating a restart_OB, PIs will get an extra 4 exposures on each LBC)

UT 03:15 - LBC crashed mid-exposure before first image completed - I received warning in Firefox GUI that an lbckill had been issued, but only people logged in are myself and OSA - IT #5697
warning email about "watchdog on Status Thread started a new instance"

UT 03:17 - Restarting M31substr6.OB

UT 03:25 - #lbcb.20150928.031804 & 032410 - Very Odd looking circular gradient - looks like reflection or light, not clear if its filter since we see stars all over

UT 03:32 - #lbcb.220150928.033022 shows same gradient, looks like its moving (these are all V-Bessell) - however, three LBC-R images look fine - they are usable

UT 03:33 - Stopping script again

UT 03:33 - Taking quick bias to see if its chip or optics related on LBC-B

UT 03:34 - LBCB bias is #033451/504 Bias looks OK

UT 03:35 - Moving back to co-pointing field to take a V, and r-sloan short exposure

UT 03:39 - copointing OB using V-Bessell (same as science) looks OK

UT 03:40 - Running DOFPIA (since been a while since focused)

UT 03:41 - Airmass 1.32 DOFPIA Complete - seeing <1"

UT 03:45 - Re-running M31substr6.OB again

UT 03:58 - So far looks OK - blue anomaly unexplained!

UT 04:12 - #040534 - LBC-B saturated - 320 sec

UT 04:15 - out of total eclipse - V-band counts now really high

UT 04:16 - I-Bessell 100sec 22k background

UT 04:20 - #041408 - saturated LBC-B

UT 04:26 - #042013 - Saturated LBC-B

UT 04:28 - M31substr6.OB partially complete

=====================================================

UT 04:30 - M31substr1_Vshorter.ob - Running DOFPIA - revised so 30sec x 2 and 80sec x8 V-Bessell (vs 60sec x1 and 320sec x2), changed I-Bessell to 50sec x2 from 100sec x 1

UT 04:34 - staring OB

UT 04:39 - 80sec V-Bessell 50k counts

UT 04:44 - OB not usable - saturated or close to it.

=====================================================

UT 04:45 - Switching from LBC to MODS as Frank program cannot be done now that we are out of totality, and Veillet Recovery isn't up until later in the night

UT 04:47 - Waking MODS

UT 05:10 - Moving to Rothberg Target IR20087-0308 ULIRG

UT 05:12 - Seeing has bloomed to ~1.2"-1.4" - airmass 1.461

UT 05:22 - Science script started

UT 05:47 - first exposure shows continuum and emission lines in Red, Blue very faint (not surprising)

UT 05:48 - Seeing is 1.3 and 1.2 avg in guider and WFS, 1.3 on DIMM airmass 1.6

UT 06:08 - dithering up 2.46" (20 pix red) to get away from bad columns, taking 3rd exposure

UT 06:31 - Moving to Feige 110

UT 06:37 - Taking spectra 5" Feige 110

UT 06:43 - Cirrus definitely visible on horizon possibly up to 45 degrees

UT 06:49 - Moving to IRF01004-2237 ULIRG Rothberg

UT 07:02 - Starting science integration

UT 08:09 - Taking a 3rd 1800sec exposure

UT 08:27 - seeing 1.1" airmass 1.7

UT 08:42 - taking 4th exposure Airmass 1.7, Seeing 1.1"

UT 09:13 - Finished ULIRG

================================================================

UT 09:14 - Moving to new target HBC340_341 (mods1r.20150928.0019-0024 ;mods1b.20150928.0011-0012 )

UT 09:27 - offsetting 0.123"left to get both stars in slit

UT 09:30 - Taking science exposure

UT 09:31 - DIMM 1.7" Guider 1.5"

UT 09:42 - Seeing is all over the place went from 0.97" in guider to 2" avg is 1.6" and 1.47" WFS, DIMM 1.7"

UT 10:12 - Thin cirrus near Moon

UT 10:15 - finished Science exposures - reconfiguring for LBCs

===================================================================

UT 10:29 - Slewing to M31subtr4_Vshorter.ob to do DOFPIA

UT 10:40 - DOFPIA converged, attempting shorter science exposures in M31substr4 field

UT 10:41 - lbcb.20150928.103821 (20k counts in V); lbcr.20150928.104036 TiO 27k counts in TiO; lbcr.2015.104614 CN 21-27k counts, 105153 I-band

UT 10:59 - modified M31subtr4_VshorterV2 to cut down on exposure times - and does last 2 dither positions

UT 11:03 - stopped now running modified script - V 40sec x16 and 15sec x4, TiO is now 150sec x2, CN 150sec x2, I-band 35sec x3 - 2 dither positions

UT 11:54 - 115258,115456 - may be affected by thin cirrus

UT 11:55 - Done with M31substr4

========================================================

UT 11:56 - Attempting Veillet recovery BZ509

UT 11:57 - slewing to target and running DOFPIA

UT 12:08 - running BZ509 - Veillet provided -0.5/mn and -0.3"/mn but NSIGUI is in "/hr -> multiplied by 60 (assuming my 5am math is correct)

UT 12:13 - put MODS to sleep

UT 12:24 - last g-sloan is at 55k counts

UT 12:26 - Only completed BZ509 - g-sloan is near saturation

=========================================================

UT 12:27 - attempting VI skyflats

UT 12:30 - Running VI skyflat scale exptime x3 5k-8k PA=0

UT 12:34 - Running again with exptime x2 - PA=0

UT 12:38 - Running Skyflat B-Bessell and CN scalex2 PA =0, CN 3k-6k

UT 12:43 - Running again, scale x 2

UT 12:45 - CN 5k-16k, last 22k

UT 12:50 - Closed

=======================================================================

-- %USERSIG{BarryRothberg - 2015-09-28}% ---++ Comments
 
Topic revision: r8 - 28 Sep 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback